# Words as Trigger Points in Social Media Discussions—A Large-Scale Case Study about UK Politics on Reddit

### v2

@@ 4.2. Animosity (Hypothesis 2)
The DiD estimator relies on the assumption of parallel trends: without the treatment, the treated and control groups would follow the same trajectory over time. While there are no formal guarantees for the assumption to hold, there are means to argue in its favour. The first argument is a theoretical one. In neither of our experimental designs, we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word. Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (semantics), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
@@
We implement event studies for each experiment to further empirically support our argument. The central idea is to model the respective outcome variable as it unfolds over time, which allows us to assess the dynamics of treatment impact before and after the intervention. Plotting the treatment effects at different time points before the treatment, the event study helps visualise whether the pre-treatment trends in the treatment and control groups are indeed parallel. If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al., 2023).
Figure 6. Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
For each of our overall 60 experiments, we divide the 30 minutes before and after the first mention of a trigger word into six distinct time intervals $\{[-30,-20[,[-20,-10[,[-10,0[,[0,10[,[10,20[,[20,30[\}${ [ - 30 , - 20 [ , [ - 20 , - 10 [ , [ - 10 , 0 [ , [ 0 , 10 [ , [ 10 , 20 [ , [ 20 , 30 [ } instead of just two. This allows us to follow, in particular, the pre-treatment trends at a higher granularity. We calculate each discussion thread’s share of (hostile) messages in the six intervals. Averaging over all these data points per period, we subtract the control group from the treatment group and calculate the difference of each observation from the first observation period.
@@
5. Conclusion
In this paper, we present the concept of trigger words and study how they affect social media conversations in UK politics. We report empirical evidence based on a large-scale study on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS and Vaccine). In short, our results show that trigger words cause more user engagement, but that this activity is marked by higher levels of animosity.
@@ 5. Conclusion
Our findings confirm the first Hypothesis (H1), which suggests more user interaction after the appearance of a trigger word. We consistently observe more posts across all trigger words and control settings. Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examine four different features that can indicate such effects. Trigger words cause more controversiality, more negative sentiment, and more anger. We also investigate the causal effect of trigger words on hate speech. Where we can, we rely on a model that is capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—. Our findings similarly confirm that trigger words play a causal role in this aspect as well.
6. Limitations
@@
We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset (>>>100 million entries). Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments. Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al., 2023).
Selecting words as trigger points and their corresponding subreddits offers room for improvement. The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community. Similarly, we only investigate a very select set of potential trigger words in the context of the UK.
However, our study had different goal: We consider our work to be a pilot study that intends to investigate whether the trigger points exist in online settings at all. Based on the empirical evidence we find, we can be confident to have found relevant empirical evidence that warrants further investigation.
7. Ethics Statement
@@
All data utilised in our research is sourced from publicly available information or collected using the official Reddit API. All information is provided in an aggregated fashion without reporting sensitive information from individual users.
Appendix A Model Categories
A.1. Emotion Categories
@@
A.2. Hate Speech Categories.
@@ Appendix B Extended Results
Figure 8 displays the average controversiality per thread before and after the introduction of the Rwanda trigger point in the Temporal and Semantic control settings.
Figure 9 displays the distribution of comments posted before and after the introduction of the Rwanda trigger point in the Temporal and Semantic control settings.
@@
Figure 7. Absolute count of posts before and after each trigger point.
@@
(b) Semantic Setting
Figure 8. Average controversiality per thread in the Temporal and Semantic settings and their respective treatment groups, for the trigger point Rwanda.
(a) Semantic Comparison
@@
Figure 9. Rwanda Treatment and Control distributions of comments in the Semantic and Temporal settings.
@@
Figure 10. Distribution of negatively charged comments in the Treatment and Control (Space) groups.
@@ Appendix C Extended Test of Parallel Trends Assumption
Figure 11. Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.

### v3

@@ 4.3 Robustness of Results
The DiD estimator relies on the assumption of parallel trends: without the treatment, the treated and control groups would follow the same trajectory over time. While there are no formal guarantees for the assumption to hold, there are means to argue in its favour. The first argument is a theoretical one. In neither of our experimental designs did we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word. Whether comparing treatment and control groups based on different Reddit fora (spaces), different words (contextual), or points in time (temporal), we suggest that without the respective trigger word, the relative difference between treatment and control group would remain the same.
@@
We implement event studies for each experiment to further empirically support our argument. The central idea is to model the respective outcome variable as it unfolds over time, which allows us to assess the dynamics of treatment impact before and after the intervention. Plotting the treatment effects at different time points before the treatment, the event study helps visualize whether the pre-treatment trends in the treatment and control groups are indeed parallel. If the trends are not parallel before the intervention, the parallel trends assumption is violated, indicating potential bias in the DID estimator (Roth et al. 2023).
[…]
@@
Placebo Treatment.
To further validate our Difference-in-Differences (DiD) estimates, we implement a placebo treatment analysis. The core logic of this approach is to check whether the estimated treatment effects could be driven by other trends earlier or later than the actual intervention.
In the placebo analysis, we systematically shift the assumed treatment time in 10-minute intervals, ranging from 30 minutes before to 30 minutes after the actual intervention, i.e., the comment with the first mention of the trigger word in a discussion thread. At each placebo point, we re-estimate all DiD models as if the treatment had occurred at that alternative time. If our results capture a genuine causal effect, we expect a pronounced treatment effect at the actual intervention time ($t=0$), and near-zero or statistically insignificant estimates at all placebo points before and after.
Figure 6 visualizes these placebo tests by plotting the precision-weighted average treatment effects and their confidence intervals across all experiments measuring animosity (see Table 5). At the actual treatment point $t=0$, which compares message content in the 30 minutes before and after the first appearance of the trigger word ([30,0[[-30,0[[ - 30 , 0 [ vs. $[0,30]$), we observe a clear and substantial effect (1.99). We then shift the treatment window rightward in 10-minute steps—first comparing [20,10[[-20,10[[ - 20 , 10 [ to $[10,40]$, and so on—until the actual intervention fully exits the analysis window at $t=30$. This generates a steady decline in estimated effects, which converge to zero, as expected if no real treatment occurred at these placebo times. Similarly, shifting the window to earlier times (e.g., comparing [60,30[[-60,-30[[ - 60 , - 30 [ to $[-30,0]$) reveals a decreasing trend in effect size, though some residual difference remains. In conclusion, given the clear peak of the effect at the timing $t=0$, we are confident in the robustness of our approach to defining the moment of the actual treatment moment.
Figure 6: Placebo treatments every 10 minutes, ranging from 30 minutes before to 30 minutes after the actual treatment point. Precision weighted average treatment effects and 95% confidence intervals across all experiments.
Figure 7: Testing the Parallel Trends Assumption with event studies. Time-wise mean of all individual event studies in purple.
5 Conclusion
This paper establishes trigger words as a concept in online debates. Since our work represents the first systematic computational study of this concept, our research design and case selection are geared to clearly identify such behaviour in the most likely cases. To do so, we performed a large-scale analysis about UK politics on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS, and Vaccine).
@@ 5 Conclusion
The results show that trigger words cause more user engagement, and this activity is marked by higher levels of animosity. This confirms the first hypothesis (H1), which suggests more user interaction after the appearance of a trigger word. Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examine four different features that can indicate such effects. Trigger words cause more controversiality, more negative sentiment, and more anger. We also investigate the causal effect of trigger words on hate speech. Where we can, we rely on a model capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—. Our findings similarly confirm that trigger words play a causal role in this aspect as well.
From this work, we can certify that trigger points exist in online communication, and we have provided a framework to measure the extent of them in terms of user engagement and animosity. Future work will analyse this phenomenon more comprehensively. We aim to broaden the scope, study how prevalent trigger points are in different countries and languages, and seek to better understand further behavioural consequences from trigger points. It is our intention to evaluate the extent to which trigger points are responsible for hateful online communication. We also strive to automate the trigger point detection process to detect the respective words in online conversations when we do not know them ex-ante.
6 Limitations
@@
We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset ($>$100 million entries). Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments. Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al. 2023).
Our study does not attempt to map how pervasive trigger points are in online debates. The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community. Similarly, we acknowledge the limitations and potential biases of our work, as we only investigate a small set of potential trigger words in the context of the UK and limit our data to comments written in English in a limited selection of subreddits. The choice of trigger words is informed by a contextual analysis of the UK politics, and the Reddit threads have been selected due to the often polarised conversations they host around such topics. However, this is the first study that analyses trigger points’ concept validity in online conversations. Based on our empirical evidence, we can be confident that we have found relevant evidence that warrants further investigation.
7 Ethics Statement
@@
All data used in our research is sourced from publicly available information or collected using the official Reddit API. It is provided in an aggregated way and without reporting sensitive information from individual users.
The automatic methods used can be abused by those with the power to repress people with opposing opinions. Hence, we offer a comprehensive study but do not share our data publicly to avoid potential misuse or commercial use. Researchers interested in replicating our experiments can contact us directly.
Appendix A Model Selection
A.1 Emotion Categories
@@
A.2 Hate Speech Categories.
@@
A.3 Validation Setting
Table 6 displays the distribution of comments in regards to emotion, sentiment and hate speech for the manually annotated dataset collected.
Emotion Hate Speech Sentiment
no emotion 29 not hate 88 negative 56
anger 26 sexism 7 neutral 37
disgust 23 racism 3 positive 7
pessimism 19 other 2
anticipation 15
sadness 9
surprise 5
optimism 5
joy 2
Table 6: Distribution of comments regarding emotion, sentiment, and hate speech categories on the manually annotated dataset.
A.4 Annotation Guidelines
Below we present the guidelines presented to annotators for the Sentiment Analysis, Emtion classification, and Hate Speech detection tasks.
Sentiment Analysis
We will show you Reddit comments that can be positive, negative, or neutral based on their sentiment.
Sentiment Categories
  • Positive comments should be labelled: Positive
  • Neutral comments should be labelled: Neutral
  • Negative comments should be labelled: Negative
Examples
  • “‘Just finished my morning workout and feeling fantastic! #HealthyLiving #FitnessGoals“‘ is a positive sentence.
  • “‘Stuck in traffic again! Late for work and already stressed out. #TrafficJam“‘ #MondayBlues is a negative sentence.
  • Factual sentences such as “‘The party was in front of the house“‘ are neutral sentences.
Notes: We have also added a comment column where you can share your thoughts, comments, or doubts. If you are unsure whether you should annotate the tweet as positive, negative or neutral, you may choose the label that suits it best and a comment about that in the comment column.
The annotation of an ironic comments depends on the context. While we acknowledge the lack of context, if you think the tweet is mocking someone (or a group of people), you should label it negative.
Emotion Classification
You will be presented with Reddit comments and asked to select the option that best describes the text. You may choose as many emotions as you believe are present — from 0 to 11 — based on the content and tone of the comment.
Emotion Categories:
  • Anger – includes annoyance, rage
  • Anticipation – includes interest, vigilance
  • Disgust – includes disinterest, dislike, loathing
  • Fear – includes apprehension, anxiety, terror
  • Joy – includes serenity, ecstasy
  • Love – includes affection
  • Optimism – includes hopefulness, confidence
  • Pessimism – includes cynicism, no confidence
  • Sadness – includes pensiveness, grief
  • Surprise – includes distraction, amazement
  • Trust – includes acceptance, liking, admiration
Notes: If no emotion is clearly present, you may leave all options unchecked. Choose all emotions that are clearly expressed or implied in the comment. If you’re unsure, use the comment column to share your reasoning or highlight ambiguity. Comments may express multiple emotions simultaneously, for example, both sadness and anger.
Hate Speech Detection
You will be presented with Reddit comments and asked to identify whether they contain hate speech, and if so, which target group is being attacked or discriminated against. Please choose only one of the following categories that best represents the nature of the content:
Hate Speech Categories:
  • Sexism: Content that targets or demeans people based on their gender (e.g., women, men, non-binary individuals) or reinforces harmful gender stereotypes. Example: “Women should stay in the kitchen where they belong.”
  • Racism: Content that expresses hostility, prejudice, or dehumanization based on race, ethnicity, or national origin. Example: “All [ethnic group] are criminals and should be deported.”
  • Disability: Content that mocks, belittles, or degrades people with physical, mental, or developmental disabilities. Example: “You’re so stupid, you must be on the spectrum.”
  • Sexual Orientation: Hate directed toward someone because of their actual or perceived sexual orientation (e.g., gay, lesbian, bisexual, etc.). Example: “Being gay is a disease that should be cured.”
  • Religion: Disparagement, dehumanization, or threats against individuals or groups based on religious beliefs or identities. Example: “All Muslims are terrorists.”
  • Other: Use this label for hate speech that targets a group not covered above, such as socioeconomic status, age, appearance, political affiliation, etc. Example: “Old people are useless and should just die already.”
  • Not Hate: Use this label if the comments does not contain hate speech. Comments that are neutral or non-discriminatory.
Notes: If you are unsure, select the label you think fits best and add a comment explaining your reasoning. A single post may mention multiple groups in such a case, select the one that is most explicit or harmful, and explain in the comment column.
@@ Appendix B Extended Results
Table 7 presents the total number of comments and threads extracted from Reddit for each keyword, along with the corresponding time periods during which the data was collected.
Table 8 presents the number of comments included in each experiment for both the treatment and control groups.
@@
Figure 9 displays the average controversiality per thread before and after the introduction of the Rwanda trigger point in the Temporal and Contextual control settings.
Figure 10 displays the distribution of comments posted before and after the introduction of the Rwanda trigger point in the Temporal and Contextual control settings.
@@
Figure 12 displays the distribution of comments conveying anger and those that do not before and after the introduction of the Vaccine trigger point in the Space control setting.
Threads Comments Year
Rwanda 0.05 1.87 2020, 2022
Brexit 0.37 11.57 2020 - 2022
NHS 0.34 6.20 2020 - 2022
Vaccine 1.42 28.11 2019, 2022
Feminism 0.94 17.79 2020 - 2022
EU 0.2 12.4 2020 - 2022
Red Cross 0.02 0.2 2020 - 2022
SubSaharan 0.08 1.5 2020, 2022
Surgery 0.5 3.9 2019, 2022
Woman 2.1 19.3 2020 - 2022
Table 7: Numbers, in millions, of threads, comments, and time period for each trigger word.
Trigger Experiment Treatment Control
Feminist space 264,495 16,503,956
contextual 163,740
Rwanda space 28,770 668,641
contextual 2,587
temp 5,896
Brexit space 688,600 10,793,036
contextual 111,524
vaccine space 348,641 7,685,095
contextual 12,986
temp 11,841
NHS space 509,458 5,275,917
contextual 1,276
Table 8: Number of comments present for each trigger in each experiment for both the treatment and control subsets.
@@
Figure 8: Absolute count of comments before and after each trigger point.
@@
(b) Contextual Setting
Figure 9: Average controversiality per thread in the Temporal and Contextual settings and their respective treatment groups, for the trigger point Rwanda.
(a) Contextual Comparison
@@
Figure 10: Rwanda Treatment and Control distributions of comments in the Contextual and Temporal settings.
@@
Figure 11: Distribution of negatively charged comments in the Treatment and Control (Space) groups.
Figure 12: Distribution of anger and non-anger comments for the treatment and control (space) groups of Vaccine trigger word.
@@ Appendix C Extended Test of Parallel Trends Assumption
Figures 13 a & b display the results of the Parallel Trends Assumption testing for all experiments as described in H1 and H2.
@@
Figure 13: Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.
@@ Appendix D Computational Resources
Appendix E Selected Reddit Comments
Below we present a Reddit discussion and how it evolves when the trigger ”feminism” appears. The comment containing the trigger is bolded.
WARNING: the following comments contain sensitive and potentially offensive language. Reader discretion is advised.
”This image though…”
”Haunting for us, isn’t it? But not for the females.”
”Well I’m sure most females would find this horrifying too.”
”You a female? Why are you on this Sub?”
”Because I’m against feminism?”
”Well, it’s better if these Posts are seen by all the Men out there. Get the taste of true Chameleon female nature.”
”Too late. I already saw it.”
”There are many women like us on this sub .”
”Yeah I’m a woman and I don’t like it. Women are not a hive mind.”
”Women are not a hive mind. You should go out and talk to one maybe.”
”My fucking eyes. Sometimes I wish I was blind.”
”When Male dies, it’s ””sometimes”” sad; when Female dies………….CRIMEEEEE!!!!
”The fucked up world we live in”